OpenStudy (anonymous):

4y/y-1 - 5y/y-2=2/y^2-3y+2 Please solve this and show all your work or explain how you did it, the book says solve by getting rid of the denominator with the LCD. THANK YOU!

OpenStudy (anonymous):

4y/y-1-5y/y-2=2/y^2-3y+2 [4y(y-2)-5y(y-1)]/y^2-3y+2=2/y^2-3y+2 -y^2-3y=2 y^2+3y+2=0 (y+1)(y+2)=0 y=-1, y=-2